### v4

@@ 2 A Design-Based Model of an Experiment
In our design-based model of an experiment, however, the data do contain information about the sample joint distribution of potential outcomes beyond the sample marginal distributions. That is, the likelihood function varies with the number of defiers within sample-level Fréchet sets. Figure 2 plots the likelihood values of each joint distribution of potential outcomes in the Fréchet set formed by the estimated marginal distributions from the second empirical example that we will discuss in Section 3. We index the joint distributions in this Fréchet set by their number of defiers along the horizontal axis. The height of the bar is the likelihood of the joint distribution in the Fréchet set with the specified number of defiers. Within this Fréchet set, the likelihood is highest at the distribution with 21 defiers, the upper Fréchet bound on defiers. As the number of defiers increases, the likelihood initially decreases, and then increases. In all empirical examples we have considered, the likelihood maintains a U-shape within the estimated Fréchet set and all other Fréchet sets, and is maximal at either the minimum or maximum number of defiers. If we normalize likelihoods in this Fréchet set to sum to one, the smallest collection of likelihoods to contain 95% of the mass excludes 9 defiers, indicated by the lighter shading.
@@
Figure 2: The Likelihood Varies with the Number of Defiers within the Estimated Fréchet Set in Johnson and Goldstein (2003), And The Smallest Collection of Likelihoods to Contain 95% of the Mass Excludes 9 Defiers (in Lighter Shading)
2.3 The Design-Based Likelihood Enables Our Proposed Maximum Likelihood Decision Rule
We consider a statistical decision problem in the style of wald1949, whereby the decision maker wishes to correctly guess the joint distribution of potential outcomes in the sample, having viewed the outcome of the experiment. In Appendix C.1, we present the formal decision problem. Statistical decision theory provides two benefits in our setting: first, decision theory is natural to apply in our design-based setting, unlike alternative criteria like consistency that depend on large-sample or asymptotic assumptions; second, statistical decision theory provides straightforward methods to quantify the gains from exploiting the variation in our likelihood over other decision rules that ignore such variation.
We propose choosing the joint distribution of potential outcomes that maximizes the design-based likelihood. Let $\widehat{{\boldsymbol{\Theta}}}({\boldsymbol{x}})$ be the set of joint distributions $\boldsymbol{\theta}$} that maximize the likelihood given the observed data $\boldsymbol{x}$}:
@@ 2.3 The Design-Based Likelihood Enables Our Proposed Maximum Likelihood Decision Rule
where $\{\bullet\}$ is the counting measure. When the likelihood is unimodal, the decision rule chooses the maximizer with probability one. We solve this optimization problem through an exhaustive grid search over the finitely many joint distributions of potential outcomes consisting of $n$ subjects.
2.4 Justifications for our Proposed Maximum Likelihood Decision Rule
Our primary motivation for proposing the maximum likelihood decision rule is that maximum likelihood rules are considered “standard” in econometric settings, and maximum likelihood estimators are familiar to practitioners. In the following subsections, we provide three additional justifications: first, that the maximum likelihood decision rule varies systematically with the data; second, that the decision rule is Bayes optimal under the appropriate conditions; and third, that the decision rule maximizes entropy, and therefore “assumes the least” (jaynes1968) about the joint distribution of potential outcomes.
2.4.1 The Maximum Likelihood Decision Rule Varies Systematically with the Data
Direct inspection of the likelihood in (2) does not reveal obvious patterns in the maximum likelihood estimate. However, by computing the maximum likelihood estimate for every possible realization of the data in fixed sample sizes, we can observe clear patterns in the estimates. Figure 3 illustrates how the maximum likelihood estimates from our proposed decision rule vary systematically with all possible realizations of the data in completely randomized experiments of 50 and 200 subjects with half in intervention. In a sample of 200 with half in intervention, there are $10,201=101\times 101$ possible realizations of the data and $1,373,301={203\choose 3}$ possible joint distributions of potential outcomes, so the visualization reflects results of grid searches over $14,009,043,501=10,201\times 1,373,301$ total elements.
Figure 3: Visualization of Our Proposed Statistical Decision Rule: How the MLE Varies with All Possible Data from Samples of 50 and 200 with Half in Intervention
In the convex lens-shaped region around the upward-sloping diagonal, MLEs only include always and never takers: the two types in the null hypothesis of the fisher1935 exact test. In samples of 50 and 200, the region in which Fisher’s exact test fails to reject the null hypothesis at the 5% level is depicted within the grey dashed line. It is larger than the region in which the MLE includes only always and never takers in both sample sizes.
The MLEs include three types in all other regions, with two regions for each combination. Even though the maximizer of the likelihood only includes two types in our stylized example from Figure 1, maximization of the likelihood sometimes requires trading off between the higher likelihood of fewer types and the higher likelihood of more balance between intervention and control within each type. If restricting the sample to two types requires too much imbalance between intervention and control, the maximizer of the likelihood includes three types. [^1] Viewing Pascal’s triangle as a representation of $N$ choose $K$ can help us to visualize the tradeoff between fewer types (moving to the next row of the triangle, increasing $N$ within each type) and more balance within each type (moving to the middle of the triangle within a row, moving $K$ closer to $N/2$ within each type). We thank Elizabeth Ananat for this point. There are no regions in which a distribution that maximizes the likelihood contains all four types. Therefore, many joint distributions of potential outcomes never maximize the likelihood for any realization of the data. For example, if a sample of 50 consisted of 18 always takers, 22 compliers, 8 defiers, and 2 never takers, the most probable realization of the data would be takeup of 20 in intervention and 13 in control, but the corresponding MLE estimate of 26 always takers, 14, compliers, 0 defiers, and 10 never takers would miss the 8 defiers entirely. We have observed MLEs that contain all four types in experiments with intervention and control arms unequal in size, although we have only observed them tied with MLEs that contain three types.
The region outlined in bold, in which the estimated average effect is positive, is of particular interest because MLEs in the shaded portion of the region include both defiers and compliers, contrary to monotonicity assumptions, as in the stylized example in Figure 1. In all even-sized samples up to 200, the MLE includes defiers when takeup is below half in control and above half in intervention, unless takeup is zero in control or full in intervention. The experiment depicted in Figure 1 falls into the region with defiers because the takeup rate is 1/3 in control and 2/3 in intervention. As shown, the MLE always includes defiers when the estimated average effect is greater than half, unless takeup is zero in control or full in intervention, and it can include defiers even when Fisher’s exact test rejects at the 5% level.
The shading demonstrates that the MLE reveals evidence beyond the average effect: holding the estimated average effect constant along upward sloping diagonal lines, MLEs can include zero or many defiers. copas1973 claims that the likelihood is always maximized at a joint distribution of potential outcomes that preserves the direct estimates of the marginal distributions of potential outcomes from equation (4) and which has either the maximal or minimal number of defiers. In practice, we often find that the likelihood is maximized at a distribution that deviates slightly from these marginal estimates.
2.4.2 The Maximum Likelihood Decision Rule is Bayes Optimal
Our proposed maximum likelihood decision rule is Bayes optimal under appropriate specifications of the decision maker’s utility function and prior. For the utility function, we specify a discrete utility function that assigns zero utility for incorrect guesses and an arbitrary amount of strictly positive utility for correct guesses; for the prior, we specify a prior that is uniform over all possible distributions of potential outcomes for $n$ subjects. Appendix C.2 reviews the formal definition and proof of Bayes optimality under these conditions. [^2] We thank Andriy Norets and Thomas Wiemann for bringing the standard conditions that establish Bayes optimality to our attention.
As we review in Appendix C.2, Bayes optimal decision rules can be derived for alternative utility functions and priors through the numerical solution of a linear programming problem, but we offer motivations for the conditions we have chosen, acknowledging their subjectivity. For the utility function, one motivation is that we can interpret it as the payoff function in a game that rewards “being correct,” as in the well-known Monty Hall game, where the contestant receives a prize only if they guess the correct door. Under this utility function, the maximum likelihood decision rule is the Maximum A Posteriori (MAP) rule, which chooses the mode of the posterior distribution. Under a uniform prior, the posterior distribution is proportional to the likelihood, and the MAP estimate equals the MLE estimate.
The desirable feature of our uniform prior is that it assigns equal weight to all distributions in the same Fréchet set, so it is “uninformative” in the sense that it does not confer any information about the joint distribution of potential outcomes beyond their marginal distributions. Furthermore, we can view our uniform prior as the result of sampling from an underlying population over which we hold an uninformative hyperprior. In sampling settings, it is common to assume a Dirichlet hyperprior over population shares. We choose a Dirichlet hyperprior over the population shares of always takers, compliers, defiers, and never takers with all concentration parameters set to one, which is equivalent to an uninformative uniform distribution for the population. Under standard assumptions on sampling from the population, the potential outcomes of subjects are independently and identically distributed (IID), and the sample distribution of potential outcomes follows a multinomial law conditional on the population distribution. Combining the Dirichlet hyperprior and the sampling process yields a Dirichlet-multinomial prior over the counts of each type in the sample. The Dirichlet-multinomial prior with concentration parameters set to one is equivalent to our uniform prior. In our stylized example in Figure 1, our uniform prior implies that the prior is the same in all three rows, so the updated posterior is proportional to the likelihood.
Regardless of the choice of prior, the ratio of the prior beliefs between any two distributions updates as long as the likelihood ratio differs from one, including within Fréchet sets. For example, consider an alternative prior that is uniform over each type for each person in the sample. This prior arises from a highly informative degenerate hyperprior that assumes the population shares of each type are equal. The resulting prior over sample distributions varies within the sample Fréchet sets, and is therefore also informative. In our stylized example in Figure 1, the resulting prior is 0.0220 for the first row, 0.0293 for the second row, and 0.0037 for the third row. The updated posterior distribution maintains curvature within the Fréchet set and updates towards the distribution in the third row that contains two defiers. To achieve a flat posterior distribution across the rows depicted in Figure 1, the prior for the first row with no defiers would have to be exactly 0.75 (=(20/8)/(20/6)) times as large as the second row with one defier and 1.5 (=(20/8)/(20/12)) times as large as the prior for the third row with two defiers.
A Bayesian perspective allows us to construct credible sets for the joint distribution of potential outcomes. In Appendix D, we describe our algorithm for constructing smallest credible sets for the joint distribution of potential outcomes. Smallest credible sets are difficult to characterize because they are multidimensional, so we also describe an algorithm for constructing smallest credible sets for the number of each type individually.
Bayes optimality ensures that our proposed decision rule does at least as well as all other decision rules under these optimality conditions. To quantify the gains from utilizing the optimal rule, we compare the Bayes expected utility obtained under the optimal rule to two alternative decision rules. The first alternative decision rule, which we refer to as the “Fréchet rule,” estimates the marginal distributions of potential outcomes directly according to (4), and then chooses randomly among the joint distributions of potential outcomes in the corresponding Fréchet set. The Fréchet rule $\widehat{{\boldsymbol{\theta}}}_{F}$ satisfies:
The second alternative decision rule, which we refer to as the “monotonicity rule,” approximates the behavior of a statistician who uses the monotonicity assumption of imbens1994 in the population to estimate the share of compliers or defiers in the sample. The rule uses the observed difference in takeup rate between the intervention and control arms $\Delta$ to approximate the share of compliers or defiers. In particular, if the difference is non-negative, the rule imposes a no-defiers monotonicity restriction, and if the difference is non-positive, it imposes a no-compliers monotonicity restriction.
$\hat{\theta}_{10}=n\cdot\max\{\Delta,0\},\quad\hat{\theta}_{01}=n\cdot\max\{-\Delta,0\}$
$=\bigg(\textbf{1}_{\{\Delta>0\}}\left[n\frac{x_{C1}}{x_{C1}+x_{C0}}\right]+\textbf{1}_{\{\Delta<0\}}\left[n\frac{x_{I1}}{x_{I1}+x_{I0}}\right],\ n\cdot\max\{\Delta,0\},\ n\cdot\max\{-\Delta,0\},$
$\qquad n-\hat{\theta}_{11}-\hat{\theta}_{01}-\hat{\theta}_{10}\bigg)$
with probability one.
We calculate the exact Bayes expected utility achieved by our optimal maximum likelihood rule and each of the two alternative rules. For a given sample size, we enumerate each of the finitely many feasible sample distributions of potential outcomes $\boldsymbol{\theta}$}. For each distribution $\boldsymbol{\theta}$}, we compute the expected utility of our decision rules by comparing the distribution of estimates they produce under each data realization $\boldsymbol{x}$} to the true distribution. We average the resulting utilities over the values of $\boldsymbol{x}$} according to the likelihood. We then average these expected utilities according to a uniform prior for $\boldsymbol{\theta}$} to get the Bayes expected utility (see Appendix C for additional detail).
Figure 4 plots the ratio of the Bayes expected utility achieved by our proposed maximum likelihood rule to the Bayes expected utility achieved by each of the two alternative rules for even sample sizes between 2 and 200. The gains from using the optimal rule over either of the alternatives increase with the sample size. In a sample size of 200, the Bayes expected utility of the proposed maximum likelihood rule is 1.79 times higher than that of the Fréchet rule, and 1.56 times higher than that of the monotonicity rule.
Figure 4: Under Optimality Conditions, For Increasing Sample Sizes, Performance of Proposed Maximum Likelihood Rule Increases Relative to Fréchet and Monotonicity Rules
2.4.3 The Maximum Likelihood Decision Rule Maximizes Entropy
We can also motivate our maximum likelihood rule through the principle of maximum entropy (jaynes1957a; jaynes1957b), which unites the theory of information with statistical physics. In Section 1.1, we discuss how differences in entropy drive variation in the likelihood such that the maximum likelihood distribution is also the maximum entropy distribution. jaynes1968 advocates for application of the principle of maximum entropy to circumvent the subjective specification of a prior. To make Bayesian decision-making more objective, one option is to choose the least informative prior. Jaynes’ alternative to make the process more objective is to choose the least informative updated distribution. The distribution that can generate the data in the greatest number of ways—the distribution that maximizes entropy, also the distribution that maximizes our design-based likelihood—is the least informative updated distribution. In the Monty Hall game, a celebrated application of Bayesian decision making, it is possible to make the same decision without specifying a Bayesian prior using a maximum likelihood decision rule and the principle of maximum entropy (wang2016).
@@ 3 Applications: Counting Defiers in Health Care
We apply our rule to two published experiments with positive, statistically significant average effects on takeup of desired health behaviors and plausible defiers. Figure 3 summarizes the results from all possible empirical applications in experiments with given sizes, illustrating empirical conditions under which our rule yields defiers. It is predictable from the data for each application and the patterns in Figure 3 that our first application yields no defiers and our second application yields defiers, which we confirm with exhaustive grid search. Our focus here is to provide examples to illustrate how applied researchers can present and interpret quantitative results and to provide substantive insights. The top of Table 1 summarizes the context and design for both applications. The statistics in the table can be calculated using our Python- and Stata-compatible dbmle package, documented at https://pypi.org/project/dbmle/ (christy2025dbmle).
In tappin2015, the sample includes 612 pregnant smokers, and the definition of “takeup” is to quit smoking, which is a desired health behavior given concerns that smoking can cause pregnancy complications and lasting adverse effects on children. The experiment uses a completely randomized design, with half of the subjects assigned to an intervention arm with a payment and the other half assigned to a control arm with usual care. Monotonicity assumptions have a long history in the context of interventions that induce pregnant women to quit smoking. In a study discussed by angrist1996, permutt1989 proposed an early monotonicity assumption requiring that women who would quit smoking in control would also quit smoking in an arm that received a multi-faceted intervention. However, literature in psychology and economics expresses concern that payments can backfire in many contexts, including on-time pickup of children from day care (gneezy2000), blood donation (mellstrom2008; lacetera2010), and vaccination (schneider2023). Defiers who would quit smoking in control but not in intervention are plausible in tappin2015 because a payment could backfire by crowding out the intrinsic motivations of pregnant women to quit smoking based on moral ideals and health goals.
The average effect of the payment relative to usual care is a statistically significant 14 percentage point increase in the smoking quit rate on a base of 8% in control. Our proposed decision rule reveals evidence beyond the average effect: our maximum likelihood estimate includes $52/612=8\%$ always takers (effect=0), $86/612=14\%$ compliers (effect=1), and $474/612=77\%$ never takers (effect=0). The weighted average effect across the three types ($8\times 0+14\times 1+77\times 0$) is equal to the estimated average effect of 14 percentage points, which is also equal to the share of compliers. We would obtain the same results if we assumed monotonicity in the sample, so our rule provides evidence in favor of monotonicity without assuming it.
We emphasize that even though our proposed decision rule provides evidence that there are no defiers in the sample, there is considerable uncertainty in our maximum likelihood estimate. As discussed in Appendix D, we report a 95% smallest credible set on the number of defiers in the sample, which we construct by reporting the collection of numbers of defiers occurring within the 95% smallest credible set on the joint distribution of potential outcomes in the sample. The smallest credible set on defiers includes up to 71 defiers, which represents 71/612=12% of the sample size.
@@
Building on the evidence provided by our decision rule, researchers could go on to make the stronger assumption of monotonicity and proceed as usual. With additional data on covariates, they could calculate average characteristics of always takers, compliers, and never takers (imbens1997; katz2001; abadie2002; abadie2003). With additional data on a second stage outcome like birth weight, they could interpret the resulting instrumental variable estimate as a local average treatment effect on compliers (imbens1994).
@@
Our second example, johnson2003, is a high-profile experiment in which a simple change from an opt-in default to an opt-out default increases organ donation almost twofold from 43% in control to 82% in intervention. Results from johnson2003 have inspired state policy changes to encourage life-saving organ donations to combat chronic organ shortages (kessler2025). However, in a recent experiment by kessler2025, in which takeup represents actually becoming an organ donor instead of hypothetically becoming an organ donor, the average effect is close to zero. Especially given the hypothetical outcome in johnson2003, both compliers and defiers are plausible. Compliers may think of the default as an endorsement by the experimenter and thus follow it regardless of what it is. Defiers may think of a default as a threat to their autonomy and thus go against it regardless of what it is.
@@
Given the plausibility of defiers in this context, a monotonicity assumption might not be reasonable, but there is no obvious alternative, so standard practice is to engage with bounds on defiers, which can be large. As shown at the bottom of Table E.1, in which we report auxiliary statistics, the smallest possible number of defiers in the sample is 0 because it is always possible for all people who take up to be always takers and all people who do not take up to be never takers, as in the null of Fisher’s exact test. The largest possible number of defiers in the sample occurs if all 11(=6150)11(=61-50) people who do not take up in intervention and all 23 people who take up in control are defiers, yielding a maximum defier share of $30\%(=(23+11)/115$). The estimated Fréchet bounds are tighter because they incorporate the estimated marginal distributions of potential outcomes in each arm using the observed takeup rates. To calculate the upper Fréchet bound from (3), we consider the bounds on the share of defiers implied by the takeup rates in intervention and control. The intervention takeup rate implies that the share of defiers in intervention and thus in the sample is at most $18\%(=100\%-82\%)$, and the control takeup rate implies that the share of defiers in control and thus in the sample is at most 43%, so the share in intervention binds, and the upper Fréchet bound is 18%.
@@
The Fréchet bounds are estimated with error, so there is even greater uncertainty in the number of defiers. The maximum number of defiers in our 95% smallest credible set is 34, which represents 30% of the sample of 115. The 95% confidence interval on the population fraction of defiers from imbens2004, incorporating the adjustment recommended by jun2023 with pretesting at the 0.01 level, also accounts for estimation of the Fréchet bounds, so we report it for comparison. It is slightly narrower, extending from 0% to 27%, but still includes the estimated Fréchet bounds. Accounting for estimation of the Fréchet bounds is empirically important. As demonstrated in Figure 2, if we do not account for estimation of the Fréchet bounds and instead construct a 95% smallest credible set on defiers within the estimated Fréchet set, variation in the likelihood allows us to exclude 9 defiers— a count in the middle of the Fréchet bounds translated into whole numbers from 0 to 21.
@@
The maximum likelihood estimate indicates that the sample includes $21/115=18\%$ defiers, which is equal to the estimated upper Fréchet bound. The maximum likelihood estimate also includes $28/115=24\%$ always takers, $66/115=57\%$ compliers, and no never takers. The maximum likelihood estimate is the same as the estimate that would be obtained under an assumption of no never takers. Under an assumption of no defiers, the estimated average effect of 39% is entirely due to compliers, who represent 39% of the sample. The maximum likelihood estimate preserves the estimated average effect ($39\%=57\%\times(1)+18\%\times(-1)$), but it tells a very different story. Compliers representing 57% of the sample are affected in one direction and defiers representing 18% of the sample are affected in the other. The total share of people affected in the maximum likelihood estimate of $75\%=(57\%+18\%)$ is almost two times as large as the number of people affected under monotonicity.
@@
In this application, evidence from the design-based model suggests a weaker alternative assumption to monotonicity—an assumption of no never takers—that can subsequently be imposed. Monotonicity can, of course, still be imposed with acknowledgment that it is a stronger assumption. While the assumption of no never takers is still strong, by the principle of maximum entropy, it “assumes the least” (jaynes1968) about the joint distribution of potential outcomes. From the design-based likelihood in (2), the assumption of no never takers can generate the data in $(28\text{ choose }15)\times(66\text{ choose }35)\times(21\text{ choose }11)=8.5\times 10^{31}$ randomized assignments representing 3.4% of the total randomized assignments ($(115\text{ choose }61)=2.5\times 10^{33}$). The assumption of no defiers can generate the data in only $(49\text{ choose }26)\times(45\text{ choose }24)\times(21\text{ choose }11)=7.8\times 10^{31}$ randomized assignments, representing 3.1% of the total randomized assignments. An ex post justification for the assumption of no never takers is that though experimental subjects might feel comfortable complying with or defying defaults imposed by the experimenter, they might feel uncomfortable revealing to the experimenter that they would not donate their organs under either default, especially since they do not know when they are asked a hypothetical question if a later question will ask them for a response under the other default. If an assumption of no never takers is not palatable, evidence for defiers could be used to motivate methods for examining effect heterogeneity, like the collection of additional data or the use of machine learning methods.

### v5

@@ 2 A Design-Based Model of an Experiment
In our design-based model of an experiment, however, the data can contain information about the sample joint distribution of potential outcomes beyond the sample marginal distributions. That is, the likelihood function can vary with the number of defiers within sample-level Fréchet sets. Figure 2 plots the likelihood values of each joint distribution of potential outcomes in the estimated Fréchet set formed by the data $(x_{I1},x_{I0},x_{C1},x_{C0})=(154,152,111,195)$, a hypothetical version of the first empirical example we discuss in Section 3. We index the joint distributions in this Fréchet set by their number of defiers along the horizontal axis. The height of the bar is the likelihood of the joint distribution in the Fréchet set with the specified number of defiers. Within this Fréchet set, the likelihood is highest at the distribution with 222 defiers, the upper Fréchet bound on defiers. As the number of defiers increases, the likelihood initially decreases, and then increases. If we normalize likelihoods in this Fréchet set to sum to one, the smallest collection of likelihoods to contain 95% of the mass excludes 105 to 116 defiers, indicated by the lighter shading. Notably, it excludes 111 defiers, the midpoint (Li and Pearl, 2019) of the Fréchet set.
@@
Figure 2: The Likelihood Varies with the Number of Defiers within the Estimated Fréchet Set, And The Smallest Collection of Likelihoods to Contain 95% of the Mass Excludes 105 to 116 Defiers (in Lighter Shading) and the Midpoint
2.3 The Design-Based Likelihood Enables Maximum Likelihood Estimation
One way to exploit the curvature in the design-based likelihood is to estimate the joint distribution of potential outcomes with maximum likelihood:
$\widehat{{\boldsymbol{\theta}}}_{MLE}({\boldsymbol{x}})={\arg\max}_{{\boldsymbol{\theta}}}\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}}).$
@@ 2.3 The Design-Based Likelihood Enables Maximum Likelihood Estimation
There are finitely many vectors of four integers that sum to the number of participants in the experiment $n$, so at least one value $\boldsymbol{\theta}$} maximizes the likelihood. We solve this optimization problem through an exhaustive grid search over the finitely many joint distributions of potential outcomes consisting of $n$ subjects.
While maximum likelihood estimation is common in practice, we also provide three justifications in Appendix B, which begins in Appendix B.1 by characterizing our estimates as the output of a statistical decision rule using the framework of statistical decision theory. First, we demonstrate that our estimates vary systematically with the data, as we show in Figure B.1 and discuss in Appendix B.2. Second, in Appendix B.3, we further justify our estimates as Bayes optimal under the appropriate conditions, including a uniform prior. A Bayesian perspective also allows us to quantify the strength of evidence by constructing credible sets. Third, in Appendix B.4, we discuss how the maximum likelihood estimate is optimal based on the principle of maximum entropy (Jaynes, 1957b, a), which “assumes the least” (Jaynes, 1968) about the joint distribution of potential outcomes conditional on the observed data.
@@ 3 Applications: Counting Defiers in Health Care
We consider two published experiments with positive, statistically significant average effects on takeup of desired health behaviors and plausible defiers. Figure B.1 summarizes the results from all possible empirical applications in experiments with given sizes, illustrating empirical conditions under which our maximum likelihood estimates yield defiers. It is predictable from the data for each application and the patterns in Figure B.1 that our first application yields no defiers and our second application yields defiers, which we confirm with exhaustive grid search. Our focus here is to provide examples to illustrate how applied researchers can present and interpret quantitative results and to provide substantive insights. The top of Table 1 summarizes the context and design for both applications. The statistics in the table can be calculated using our Python- and Stata-compatible dbmle package, documented at https://pypi.org/project/dbmle/ (Christy, Kowalski, and Zhang, 2025).
@@
As in our stylized example, in Tappin et al. (2015), the definition of “takeup” is to quit smoking, the intervention involves payment, and the design is completely randomized with half assigned to intervention. The sample includes 612 pregnant smokers. Monotonicity assumptions have a long history in the context of interventions that induce pregnant women to quit smoking. In a study prominently discussed by Angrist, Imbens, and Rubin (1996), Permutt and Hebel (1989) propose an early monotonicity assumption requiring that women who would quit smoking in control would also quit smoking in an arm that received a multifaceted intervention. However, the intervention in Tappin et al. (2015) involves payment, and as the psychologist friend warned in the introduction, literature since the monotonicity assumption was proposed expresses concern that payments can backfire in many contexts, including on-time pickup of children from day care (Gneezy and Rustichini, 2000), blood donation (Mellström and Johannesson, 2008; Lacetera and Macis, 2010), and vaccination (Schneider, Campos-Mercade, Meier, Pope, Wengström, and Meier, 2023). Defiers who would quit smoking in control but not in intervention are plausible in Tappin et al. (2015) because a payment could backfire by crowding out the intrinsic motivations of pregnant women to quit smoking based on moral ideals and health goals.
The average effect of the payment relative to usual care is a statistically significant 14 percentage point increase in the smoking quit rate on a base of 8% in control. Our maximum likelihood estimates reveal evidence beyond the average effect: $52/612=8\%$ always takers (effect=0), $86/612=14\%$ compliers (effect=1), $0/612=0\%$ defiers (effect=-1), and $474/612=77\%$ never takers (effect=0). The weighted average effect across the four types ($8\times 0+14\times 1+0\times(-1)+77\times 0$) is equal to the estimated average effect of 14 percentage points, which is also equal to the share of compliers. We would obtain the same results if we assumed monotonicity in the sample, so our estimates provide evidence in favor of monotonicity without assuming it.
Returning to the concerns of the doctor in the introduction, the maximum likelihood estimates are reassuring. No patients would be harmed by payment—the doctor’s greatest concern. The 8% who are always takers represent a modest waste, although the doctor is not too disappointed about giving payments to pregnant women who quit smoking. The 77% of patients who are never takers cannot be reached by the intervention, which is disappointing but not costly given that payment is contingent on quitting. The 14% who are compliers are exactly the patients the doctor hoped to help.
Our design-based perspective affects the interpretation of our results. The doctor in the introduction cares about the patients in the sample, and our design-based likelihood is informative because it conveys information about the sample at hand. However, if the doctor were interested in the effects of a payment policy in the population from which the sample was drawn, the sampling-based likelihood we derive in Appendix A.2 would reveal no information about defiers in the population beyond the Fréchet bounds—at least, not without a departure from standard sampling assumptions, which we explore in Appendix A.3.
We emphasize that even though our maximum likelihood estimates provide evidence that there are no defiers in the sample, considerable uncertainty remains. As discussed in Appendix B.3, we compute a 95% smallest credible set on the joint distribution of potential outcomes, and we report the ranges for the numbers of always takers, compliers, defiers, and never takers in that set in Table 1. The 95% smallest credible set includes distributions with 0 to 71 defiers, representing 0% to 71/612=12% of the sample. The magnitudes of the values in this 95% smallest credible set help us to highlight that evidence on defiers is weak when the estimated average effect is positive.
@@
Building on the evidence here, researchers could go on to make the stronger assumption of monotonicity and proceed as usual. With additional data on covariates, they could calculate average characteristics of always takers, compliers, and never takers (Imbens and Rubin, 1997; Katz, Kling, Liebman, and others, 2001; Abadie, 2002, 2003). With additional data on a second stage outcome like birth weight, they could interpret the resulting instrumental variable estimate as a local average treatment effect on compliers (Imbens and Angrist, 1994).
Note that the joint distribution of potential outcomes can change dramatically even when the average effect does not. If the control takeup rate in Tappin et al. (2015) were $111/306=36\%$ instead of 8%, with the same average effect of 14 percentage points (takeup among an additional 43 people in intervention), the maximum likelihood estimates would include 36% defiers rather than zero. Note that these are the same data used for Figure 2 at the end of Section 2.2. Our second application illustrates the estimation of a positive count of defiers with real data.
@@
Our second application, Johnson and Goldstein (2003), is a high-profile experiment in which a simple change from an opt-in default to an opt-out default increases organ donation almost twofold from 43% in control to 82% in intervention. Results from Johnson and Goldstein (2003) have inspired state policy changes to encourage life-saving organ donations to combat chronic organ shortages (Kessler and Roth, 2025). However, in a recent experiment by Kessler and Roth (2025), in which takeup represents actually becoming an organ donor instead of hypothetically becoming an organ donor, the average effect is close to zero. Especially given the hypothetical outcome in Johnson and Goldstein (2003), both compliers and defiers are plausible. Compliers may think of the default as an endorsement by the experimenter and thus follow it regardless of what it is. Defiers may think of a default as a threat to their autonomy and thus go against it regardless of what it is. For this application, the population from which the sample is drawn has clear policy relevance, so we begin by engaging with standard sampling-based results.
@@
Given the plausibility of defiers in this context, a monotonicity assumption might not be reasonable, but there is no obvious alternative, so standard practice is to engage with bounds on defiers, which can be large. As shown at the bottom of Table C.3, in which we report auxiliary statistics, the smallest possible number of defiers in the sample is 0 because it is always possible for all people who take up to be always takers and all people who do not take up to be never takers, as in the null of Fisher’s exact test. The largest possible number of defiers in the sample occurs if all 11(=6150)11(=61-50) people who do not take up in intervention and all 23 people who take up in control are defiers, yielding a maximum defier share of $30\%(=(23+11)/115$). The estimated Fréchet bounds are tighter because they incorporate the estimated marginal distributions of potential outcomes in each arm using the observed takeup rates. To calculate the upper Fréchet bound from (3), we consider the bounds on the share of defiers implied by the takeup rates in intervention and control. The intervention takeup rate implies that the share of defiers in intervention and thus in the sample is at most $18\%(=100\%-82\%)$, and the control takeup rate implies that the share of defiers in control and thus in the sample is at most 43%, so the share in intervention binds, and the upper Fréchet bound is 18%.
@@
The Fréchet bounds are estimated with error, so there is even greater uncertainty in the number of defiers. The number of defiers in our 95% smallest credible set ranges from 0 to 34, which represents 0% to 30% of the sample of 115. The 95% confidence interval on the population fraction of defiers from Imbens and Manski (2004), incorporating the adjustment recommended by Jun and Lee (2023) with pretesting at the 0.01 level, also accounts for estimation of the Fréchet bounds, so we report it for comparison in Table C.3. It is slightly narrower, extending from 0% to 27%, but still includes the estimated Fréchet bounds. Accounting for estimation of the Fréchet bounds is empirically important. As demonstrated in Table C.3, if we do not account for estimation of the Fréchet bounds and instead construct a 95% smallest credible set within the estimated Fréchet set, variation in the likelihood allows us to exclude 9 defiers—a count in the middle of the Fréchet bounds translated into whole numbers from 0 to 21.
@@
The maximum likelihood estimates indicate that the sample includes $21/115=18\%$ defiers, which is equal to the estimated upper Fréchet bound. The maximum likelihood estimates also include $28/115=24\%$ always takers, $66/115=57\%$ compliers, and no never takers. The maximum likelihood estimates are the same as the estimate that would be obtained under an assumption of no never takers. Under an assumption of no defiers, the estimated average effect of 39% is entirely due to compliers, who represent 39% of the sample. The maximum likelihood estimates preserve the estimated average effect ($39\%=57\%\times(1)+18\%\times(-1)$), but they tell a very different story. Compliers representing 57% of the sample are affected in one direction and defiers representing 18% of the sample are affected in the other. The total share of people affected of $75\%=(57\%+18\%)$ is almost two times as large as the number of people affected under monotonicity.
@@
In this application, evidence from the design-based model suggests a weaker alternative assumption to monotonicity—an assumption of no never takers—that can subsequently be imposed. Monotonicity can, of course, still be imposed with acknowledgment that it is a stronger assumption. While the assumption of no never takers is still strong, by the principle of maximum entropy, it “assumes the least” (Jaynes, 1968) about the joint distribution of potential outcomes. From the design-based likelihood in (9), the assumption of no never takers can generate the data in at most $(28\text{ choose }50-(66-(54-23)))\times(66\text{ choose }66-(54-23))\times(21\text{ choose }(61-50))=8.5\times 10^{31}$ randomized assignments representing 3.4% of the total randomized assignments ($(115\text{ choose }61)=2.5\times 10^{33}$). The assumption of no defiers can generate the data in at most $(49\text{ choose }26)\times(45\text{ choose }24)\times(21\text{ choose }11)=7.8\times 10^{31}$ randomized assignments, representing 3.1% of the total randomized assignments. An ex post justification for the assumption of no never takers is that though experimental subjects might feel comfortable complying with or defying defaults imposed by the experimenter, they might feel uncomfortable revealing to the experimenter that they would not donate their organs under either default, especially since they do not know when they are asked a hypothetical question if a later question will ask them for a response under the other default. If an assumption of no never takers is not palatable, evidence for defiers could be used to motivate methods for examining effect heterogeneity, like the collection of additional data or the use of machine learning methods.